OpenStudy (anonymous):

A bag contains 8 blue, 3 red, and 6 white chips. Only red chips are added to the bag. How many red chips must be added to the bag for the probability of drawing a red chip to be 1/3?

OpenStudy (anonymous):

Red chips at the moment = 3 Total chips at the moment = 17 Say we need to add x red chips to make the probablity of drawing a red chip = 1/3 Thus, total number of red chips = 3 + x Total number of chips = 17 + x Thus, Probability of drawing a red chip = (3+x/ 17 + x) = 1/3 3(3+x) = 17+ x. Solve for x. Hope this helps